A thin hollow sphere of mass ‘m’ is completely filled with a liquid of mass ‘m’.  When the sphere rolls with a velocity ‘v’, kinetic energy of the system is (neglect friction):


 



The maximum tension a rope can withstand is 60 kg wt.  The ration of maximum acceleration with which two boys of masses 20 kg and 30 kg can climb up the rope ate the same time is:

Grade:11

2 Answers

MADDALA NAGENDRAKUMAR
4 years ago
For 20 kg body FBD equation is
W1=T+m1(g+a1)
200=60+20(10+a1)
200-200=60+20a1
a1=60/20=3
For 30kg body FBD Equation is
W2=T+m2(g+a2)
300=60+30(10+a2)
300-300=60+30a2
a2=60/30=2
a1:a2=3:2
So finally the ratio between a1 and a2 is
(a1:a2
MADDALA NAGENDRAKUMAR
4 years ago
FOR 20KG BOY FBD EQUATION IS
M1a1=T-W1
20a1=600-200
20a1=400
a1=20
FOR 30KG BOY FBD EQUATION IS
30a2=T-W2
30a2=600-300
30a2=300
a2=10
the ratio of a1 and a2 is
a1:a2=20:10
a1:a2=2:1